OpenStudy (anonymous):

Which equation would you use to solve the following word problem? Sasha's mom decided to get her a cell phone for her birthday. Cell Plus has a plan that costs $23.95 per month plus an additional $0.05 per minute. Cell Best has a plan that costs $28.95 per month plus an additional $0.03 per minute. How many minutes can Sasha talk and have the same cost each month? 0.08m = 5.00 23.95 + 0.05m = 28.98 28.95 + 0.03m = 24.00 23.95 + 0.05m = 28.95 + 0.03m

OpenStudy (anonymous):

23.95 + 0.05m = 28.95 + 0.03m

OpenStudy (anonymous):

I'm guessing it would be the last one because it wants you to compare the two? >_< I don't know it's a bit confusing because it gives you two different problems 23.95+0.05m=x and 28.95+0.03m=x But I'm thinking it's asking how many minutes would she need to talk on Cell plus, to get the same cost per month as Cell Best, sooo 23.95+0.05m= 28.98 seems like the best answer

OpenStudy (anonymous):

last one O:

OpenStudy (anonymous):

Thanks both of you

OpenStudy (anonymous):

you gonna set 2 different equations in order to find the intersection

OpenStudy (anonymous):

Last one what?

OpenStudy (anonymous):

Ohh get rid of the first sentence in mine, i thought it was the last one at first until I thought about it more

OpenStudy (anonymous):

once you find where then mean the x will be the amount of minutes one has to talk so that the two plan could cost the same

OpenStudy (anonymous):

meet* not mean Dx

OpenStudy (anonymous):

So whats the answer lol @Nemo1223

OpenStudy (anonymous):

23.95 + 0.05m = 28.95 + 0.03m put in it a graphing calculator and find intersection to find for m

OpenStudy (anonymous):

Okay so its d